# Explainable Multi-modal Time Series Prediction with LLM-in-the-Loop

### v1

@@ 1 Introduction
In the modern big-data era, time series analysis has become indispensable for understanding real-world system behaviors and guiding downstream decision-making tasks across numerous domains, including healthcare, traffic, finance, and weather (Jin et al., 2018; Guo et al., 2019; Zhang et al., 2017; Qin et al., 2017). Although deep learning models have demonstrated success in capturing complex temporal dependencies (Nie et al., 2023; Deng & Hooi, 2021; Zhang et al., 2022; Liu et al., 2024d), real-world time series are frequently influenced by external information beyond purely temporal factors. Such additional context, which may come from textual narratives (e.g., finance news (Dong et al., 2024) or medical reports (King et al., 2023)), can offer critical insights for more accurate forecasting and explainability.
@@
Recent multi-modal approaches for time series have shown promise by integrating rich contextual signals from disparate data sources—such as textual descriptions—to improve performance on tasks ranging from forecasting and classification to imputation and retrieval (Ekambaram et al., 2020; Niu et al., 2023; Lee et al., 2024; Xing & He, 2023; Moroto et al., 2024; Zhao et al., 2022; Bamford et al., 2023). While these approaches utilize supplementary data to enhance predictive accuracy, they often lack explicit mechanisms to systematically reason and explain about why or how contextual signals affect outcomes. This gap in interpretability poses significant barriers for high-stakes applications such as finance or healthcare, where trust and transparency are paramount.
@@
Meanwhile, Large Language Models (LLMs) (Achiam et al., 2023; Team et al., 2023; Touvron et al., ) have risen to prominence for their remarkable ability to process and reason over textual data across domains, enabling tasks like sentiment analysis, question answering, and content generation in zero- and few-shot settings (Zhang et al., 2024; Kamalloo et al., 2023; Wang et al., 2024c). Their encoded domain knowledge makes them natural candidates for supporting multi-modal time series analyses, where textual context (e.g., news or expert notes) plays a vital role (Liu et al., ; Nie et al., 2024; Koa et al., 2024; Wang et al., 2023; Shi et al., 2024; Yu et al., 2023; Singhal et al., 2023).
@@
Figure 1: An overview of the TimeXL workflow. A prototype-based explainable encoder first produces predictions and case-based rationales for both time series and text. The prediction LLM refines forecasts based on these rationales (Step 1). A reflection LLM then critiques the output against ground truth (Step 2), providing feedback to detect textual noise (Step 3). Finally, a refinement LLM updates the text accordingly, triggering encoder retraining for improved accuracy and explanations (Step 4).
@@
Motivated by these observations, we introduce TimeXL, a novel framework that adopts a closed-loop workflow of prediction, critique (reflect), and refinement, and unifies a prototype-driven time series encoder with LLM-based reasoning to deliver both accurate and interpretable multi-modal forecasting (Figure 1). Our approach first employs a multi-modal prototype-based encoder to generate preliminary time series predictions alongside human-readable explanations, leveraging case-based reasoning (Kolodner, 1992; Ming et al., 2019; Ni et al., 2021; Jiang et al., 2023) from both the temporal and textual modalities. These explanations not only justify the encoder’s predictions but also serve as auxiliary signals to guide an LLM-powered component that further refines the forecasts and contextual rationales.
@@
  • Experiments on four real-world benchmarks show that TimeXL consistently outperforms baselines, achieving up to a 8.9% improvement in AUC while providing faithful, human-centric multi-modal explanations.
Overall, TimeXL opens new avenues for explainable multi-modal time series analysis by coupling prototype-based inference with LLM-driven reasoning.
@@ 2 Related Work
2.1 Multi-modal Time Series Analysis
In recent years, multi-modal time series analysis has gained significant traction in diverse domains such as finance, healthcare, environmental sciences, and industry (Ekambaram et al., 2020; Skenderi et al., 2024; Niu et al., 2023; Yang & Wu, 2021; Zhao et al., 2022; Xing & He, 2023). Multiple approaches have been proposed to model interactions across different modalities for various tasks. For instance, (Lee et al., 2024) introduces a multi-modal augmentation framework for few-shot time series forecasting, which fuses time series and textual representations both at the sample and feature levels using attention. Furthermore, (Bamford et al., 2023) aligns multi-modal time series within a shared latent space of deep encoders and retrieves specific sequences based on textual queries. In addition, (Zheng et al., 2024) performs causal structure learning to uncover root causes in multi-modal time series by separating modality-invariant and modality-specific components via contrastive learning. Most recently, (Liu et al., 2024a) established a multi-modal forecasting benchmark with baselines, and demonstrating performance improvements through the incorporation of a new modality. Although these techniques have advanced predictive performance by leveraging cross-modality interactions, they tend to focus primarily on improving numerical accuracy. The deeper reasoning behind how or why the textual or other contextual signals influence time series outcomes remains underexplored.
2.2 Time Series Explanation
@@
Recent studies have explored diverse paradigms for time series interpretability. Gradient-based and perturbation-based “saliency” methods, for example, highlight important features at different time steps (Ismail et al., 2020; Tonekaboni et al., 2020), while other works explicitly incorporate temporal structures into models and objectives (Leung et al., ; Crabbé & Van Der Schaar, 2021). Surrogate approaches also offer global or local explanations, such as applying Shapley values to time series (Bento et al., 2021), enforcing model consistency via self-supervised objectives (Queen et al., 2024), or using information-theoretic strategies for coherent explanations (Liu et al., 2024f). In contrast to saliency or surrogate-based explanations, we adopt a case-based reasoning paradigm (Kolodner, 1992; Ming et al., 2019; Ni et al., 2021; Jiang et al., 2023), which end-to-end generates predictions and built-in explanations from learned prototypes. Our work extends this approach to multi-modal time series by producing human-readable reasoning artifacts for both the temporal and contextual modalities.
2.3 LLMs for Time Series Analysis
@@
The rapid development of Large Language Models (LLMs) (Achiam et al., 2023; Team et al., 2023; Touvron et al., ) has begun to inspire new directions in time series research (Jiang et al., 2024; Liang et al., 2024). Many existing techniques fine-tune pre-trained LLMs on time series tasks, achieving state-of-the-art results in forecasting, classification, and beyond (Zhou et al., 2023; Bian et al., ; Ansari et al., 2024). Often, textual data—such as domain instructions, metadata, or dataset summaries—are encoded as prefix embeddings to enrich time series representations (Jin et al., 2024; Liu et al., 2024b; Jia et al., 2024; Liu et al., 2025). These techniques also contribute to the emergence of time series foundation models (Ansari et al., 2024; Das et al., 2023; Woo et al., 2024; Liu et al., 2024e; Wang et al., 2024a). An alternative line of research leverages the zero-shot or few-shot reasoning capabilities of LLMs. These methods directly prompt pre-trained language models with text-converted time series (Xue & Salim, 2023) or context-laden prompts representing domain knowledge (Wang et al., 2023; Yu et al., 2023; Singhal et al., 2023), often yielding surprisingly strong performance in real-world scenarios. Furthermore, LLMs can act as knowledge inference modules, synthesizing high-level patterns or explanations that augment standard time series pipelines (Chen et al., 2023b; Shi et al., 2024; Lee et al., 2025; Wang et al., 2024b).
@@ 3 Methodology
Figure 2: The overview of TimeXL (left), and multi-modal prototype-based encoder (right).
@@ 3.1 Problem Statement
In this paper, we consider a multi-modal time series prediction problem. Each instance is represented by the multi-modal input $(\boldsymbol{x},\boldsymbol{s})$, where $\boldsymbol{x}=(x_{1},x_{2},\cdots,x_{T})\in\mathbb{R}^{N\times T}$ denotes time series data with $N$ variables and $T$ historical time steps, and $\boldsymbol{s}$ denotes the corresponding text data describing the real-world context. The text data $\boldsymbol{s}$ can be further divided into $L$ meaningful segments. Based on the historical time series and textual context, our objective is to predict the future outcome $\boldsymbol{y}$, either as a discrete value for classification tasks, or as a continuous value for regression tasks. In this paper, we mainly consider a classification task while we provide a demonstration of the regression task in Appendix F. There are four major components in the proposed TimeXL framework, a multi-modal prototype encoder $\mathcal{M}_{\mathrm{enc}}$ that provides initial prediction and case-based explanation, a prediction LLM $\mathcal{M}_{\mathrm{pred}}$ that provides prediction based on the understanding of context with explanation, a reflection LLM $\mathcal{M}_{\mathrm{refl}}$ that generates feedback, and a refinement LLM $\mathcal{M}_{\mathrm{refine}}$ that refines the textual context based on the feedback. Below, we introduce each component and how they synergize toward better prediction and explanation.
@@ 3.2.1 Multi-modal Sequence Modeling with Prototypes
Sequence Encoder. To capture both temporal and semantic dependencies, we adopt separate encoders for time series ($\mathcal{E}_{\theta}$) and text ($\mathcal{E}_{\phi}$). For $\boldsymbol{x}\in\mathbb{R}^{N\times T}$, the time series encoder $\mathcal{E}_{\theta}$ maps the entire sequence into one or multiple representations, which serve as candidates for prototype learning. Simultaneously, the text input $\boldsymbol{s}$ is first transformed by a frozen pre-trained language model, $\mathrm{PLM}$ (e.g., BERT(Kenton & Toutanova, 2019) or Sentence-BERT(Reimers & Gurevych, 2019)), to produce embeddings $\boldsymbol{e}_{s}\in\mathbb{R}^{d_{s}\times L}$. These embeddings are then processed by a separate encoder $\mathcal{E}_{\phi}$ to extract meaningful text features. It is worth noting that the choice of $\mathcal{E}_{\theta}$ and $\mathcal{E}_{\phi}$ also affects the granularity of explanations. As we will introduce shortly, the prototypes are learned based on sequence representations and are associated with the counterparts in the input space, where the correspondences are determined by the encoders. In this paper, we choose convolution-based encoders for both modalities to capture the fine-grained sub-sequence (i.e., segment) patterns:
$\boldsymbol{Z}_{\mathrm{time}}$ $=\bigl{(}\boldsymbol{z}_{1},\dots,\boldsymbol{z}_{T-w+1}\bigr{)}\;=\;\mathcal{E}_{\theta}(\boldsymbol{x}),$ (1)
$\boldsymbol{Z}_{\mathrm{text}}$ $=\bigl{(}\boldsymbol{z}^{\prime}_{1},\dots,\boldsymbol{z}^{\prime}_{L-w^{\prime}+1}\bigr{)}\;=\;\mathcal{E}_{\phi}(\boldsymbol{e}_{s}),$ (2)
@@
Prototype Allocation. To establish interpretability, we learn a set of time series prototypes and text prototypes for each class $c\in\{1,\dots,C\}$. Specifically, we introduce:
$\boldsymbol{P}^{(c)}_{\mathrm{time}}\in\mathbb{R}^{k\times h},\quad\boldsymbol{P}^{(c)}_{\mathrm{text}}\in\mathbb{R}^{k^{\prime}\times h^{\prime}},$
so that each prototype $\boldsymbol{p}_{i}^{(c)}\in\mathbb{R}^{h}$ (time series) or $\boldsymbol{p}_{i}^{\prime(c)}\in\mathbb{R}^{h^{\prime}}$ (text) resides in the same feature space as the relevant encoder outputs. For an input sequence, we measure the similarity between each prototype and the most relevant segment in the corresponding modality:
$\mathrm{Sim}_{i}^{(c)}=\operatorname{max}(\mathrm{Sim}_{i,1}^{(c)},\cdots,\mathrm{Sim}_{i,T-w+1}^{(c)})$
$\text{where}\;\;\mathrm{Sim}_{i,j}^{(c)}=\exp\left(-\left\|\boldsymbol{p}^{(c)}_{i}-\boldsymbol{z}_{j}\right\|_{2}^{2}\right)\in[0,1]$ (3)
@@ 3.2.2 Learning Prototypes toward Better Explanation
$\mathcal{L}_{c}=\sum_{\boldsymbol{z}_{j}\in\boldsymbol{Z}_{(\cdot)}}\min_{\boldsymbol{p}_{i}\in\boldsymbol{P}_{(\cdot)}}\left\|\boldsymbol{z}_{j}-\boldsymbol{p}_{i}\right\|_{2}^{2},$ (5)
$\mathcal{L}_{e}=\sum_{\boldsymbol{p}_{i}\in\boldsymbol{P}_{(\cdot)}}\min_{\boldsymbol{z}_{j}\in\boldsymbol{Z}_{(\cdot)}}\left\|\boldsymbol{p}_{i}-\boldsymbol{z}_{j}\right\|_{2}^{2}$
@@ 3.3.1 Model Synergy for Augmented Prediction
Prediction with Enriched Contexts. The prediction LLM agent $\mathcal{M}_{\mathrm{pred}}$ generates predictions based on the input text $\boldsymbol{s}$. To improve prediction accuracy, the encoder $\mathcal{M}_{\mathrm{enc}}$ supplements $\boldsymbol{s}$ with case-based explanations. Specifically, $\mathcal{M}_{\mathrm{enc}}$ selects the $\omega$ prototypes that exhibit the highest relevance to any of the textual segments within $\boldsymbol{s}$. Relevance is determined by the similarity scores used in Equation 3.2.1. These selected prototypes are then added to the input prompt of $\mathcal{M}_{\mathrm{pred}}$ as explanations, providing richer real-world context and leading to more accurate predictions. The $\omega$ prototype-segment pairs, which construct the explanation $\mathbf{expl}_{s}$ of the input text $\boldsymbol{s}$, are retrieved as follows:
$\mathbf{expl}_{s}=\left\{\left(\boldsymbol{p}_{i}^{(c)},\boldsymbol{s}_{j}\right):(i,j,c)\in\operatorname{Top-\omega}(\mathrm{Sim}_{\mathrm{text}})\right\}$
where $\text{where }\operatorname{Top-\omega}(\mathrm{Sim}_{\mathrm{text}})=\operatorname{argTop-\omega}_{(i,j,c)}\left(\mathrm{Sim’}_{i,j}^{(c)}\right).$
Note that $i,j,c$ denotes the prototype index, segment index, and class index, respectively. As $\mathbf{expl}_{\boldsymbol{s}}$ can contain relevant contextual guidance across multiple classes, it augments the input space and removes semantic ambiguity for prediction agent $\mathcal{M}_{\mathrm{pred}}$. Therefore, the prediction is drawn as $\boldsymbol{\hat{y}}_{\mathrm{LLM}}=\mathcal{M}_{\mathrm{pred}}(\boldsymbol{s},\mathbf{expl}_{\boldsymbol{s}})$. The prompt for querying the prediction agent $\mathcal{M}_{\mathrm{pred}}$ is provided in Appendix D, Figure 12.
Fused Predictions. We compile the final prediction based on a fusion of both the multi-modal encoder $\mathcal{M}_{\mathrm{enc}}$ and prediction LLM $\mathcal{M}_{\mathrm{pred}}$. Specifically, we linearly combine the continuous prediction probabilities $\boldsymbol{\hat{y}}_{\mathrm{enc}}$ and discrete prediction $\boldsymbol{\hat{y}}_{\mathrm{LLM}}$: $\boldsymbol{\hat{y}}=\alpha\boldsymbol{\hat{y}}_{\mathrm{enc}}+(1-\alpha)\boldsymbol{\hat{y}}_{\mathrm{LLM}}$, where $\alpha\in[0,1]$ is the hyperparameter selected from validation data. The encoder $\mathcal{M}_{\mathrm{enc}}$ and prediction agent $\mathcal{M}_{\mathrm{pred}}$ enhance each other based on their unique strengths. The $\mathcal{M}_{\mathrm{enc}}$ is fine-tuned based on explicit supervised signals, ensuring accuracy in capturing temporal and contextual dependencies of multi-modal time series. On the other hand, $\mathcal{M}_{\mathrm{pred}}$ contributes deep semantic understanding drawn from extensive text corpora. By fusing predictions from two distinct perspectives, we achieve a synergistic augmentation toward more accurate and comprehensive predictions for complex multi-modal time series.

### v2

@@ 1 Introduction
In the modern big-data era, time series analysis has become indispensable for understanding real-world system behaviors and guiding downstream decision-making tasks across numerous domains, including healthcare, traffic, finance, and weather [1, 2, 3, 4]. Although deep learning models have demonstrated success in capturing complex temporal dependencies [5, 6, 7, 8], real-world time series are frequently influenced by external information beyond purely temporal factors. Such additional context, which may come from textual narratives (e.g., finance news [9] or medical reports [10]), can offer critical insights for more accurate forecasting and explainability.
@@
Recent multi-modal approaches for time series have shown promise by integrating rich contextual signals from disparate data sources—such as textual descriptions—to improve performance on tasks ranging from forecasting and classification to imputation and retrieval [11, 12, 13, 14, 15, 16, 17]. While these approaches utilize supplementary data to enhance predictive accuracy, they often lack explicit mechanisms to systematically reason and explain about why or how contextual signals affect outcomes. This gap in interpretability poses significant barriers for high-stakes applications such as finance or healthcare, where trust and transparency are paramount.
@@
Meanwhile, Large Language Models (LLMs) [18, 19, 20] have risen to prominence for their remarkable ability to process and reason over textual data across domains, enabling tasks like sentiment analysis, question answering, and content generation in zero- and few-shot settings [21, 22, 23]. Recent advancements in agentic designs further augment LLM capabilities, facilitating structured reasoning and iterative decision-making for context-rich scenarios [24, 25, 26]. Their encoded domain knowledge makes them natural candidates for supporting real-world multi-modal time series analyses, where textual context (e.g., news or expert notes) plays a vital role [27, 28, 29, 30, 31, 32, 33].
@@
Figure 1: An overview of the TimeXL workflow. A prototype-based explainable encoder first produces predictions and case-based rationales for both time series and text. The prediction LLM refines forecasts based on these rationales (Step 1). A reflection LLM then critiques the output against ground truth (Step 2), providing feedback to detect textual noise (Step 3). Finally, a refinement LLM updates the text accordingly, triggering encoder retraining for improved accuracy and explanations (Step 4). Note that the predictions from the encoder and the LLM are also fused to enhance overall accuracy.
@@
Motivated by these observations, we introduce TimeXL, a novel framework that adopts a closed-loop workflow of prediction, critique (reflect), and refinement, and unifies a prototype-driven time series encoder with LLM-based reasoning to deliver both accurate and interpretable multi-modal forecasting (Figure 1). Our approach first employs a multi-modal prototype-based encoder to generate preliminary time series predictions alongside human-readable explanations, leveraging case-based reasoning [34, 35, 36, 37] from both the temporal and textual modalities. These explanations not only justify the encoder’s predictions but also serve as auxiliary signals to guide an LLM-powered component that further refines the forecasts and contextual rationales.
@@
  • Experiments on four real-world benchmarks show that TimeXL consistently outperforms baselines, achieving up to a 8.9% improvement in AUC while providing faithful, human-centric multi-modal explanations. Overall, TimeXL opens new avenues for explainable multi-modal time series analysis by coupling prototype-based inference with LLM-driven reasoning.
@@ 2 Related Work
Our work is primarily related to three key areas of time series research: multi-modality, explanation, and the use of LLMs. A more detailed discussion of related methods is provided in Appendix G.
Multi-modal Time Series Analysis. In recent years, multi-modal time series analysis has gained increasing attention in diverse domains such as finance, healthcare, environmental sciences, and industry [11, 38, 12, 39, 16, 14]. Multiple approaches have been proposed to model interactions across different modalities, including fusion and alignment through concatenation [40, 41], attention mechanism [13, 42, 43], gating functions [44], and contrastive learning [45, 46]. These methods facilitate tasks beyond forecasting and classification, such as retrieval [17], imputation [16], and causal discovery [45]. Recent research efforts have also advanced the field through comprehensive benchmarks and studies [47, 48, 49, 50, 51], highlighting the incorporation of new modalities to boost task performance. Meanwhile, others explore transforming time series to other modalities [52, 53, 54, 55] or generating synthetic ones [44, 56] for effective analysis. Nevertheless, these studies primarily focus on improving numerical performance while the deeper and tractable rationale behind why or how the textual or other contextual signals influence time series outcomes remains underexplored.
@@
Time Series Explanation. Recent studies have explored diverse paradigms for time series interpretability. Gradient-based and perturbation-based saliency methods, for example, highlight important features at different time steps [57, 58], where some works explicitly incorporate temporal structures into models and objectives [59, 60]. Surrogate approaches also offer global or local explanations, such as applying Shapley values to time series [61], enforcing model consistency via self-supervised objectives [62], or using information-theoretic strategies for coherent explanations [63]. In contrast to saliency or surrogate-based explanations, we adopt a case-based reasoning paradigm [34, 35, 36, 37], which end-to-end generates predictions and built-in explanations from learned prototypes. Our work extends this approach to multi-modal time series by producing human-readable reasoning artifacts for both the temporal and contextual modalities.
@@
LLMs for Time Series Analysis. The rapid development of LLMs [18, 19, 20] has begun to inspire new directions in time series research [64, 65]. Many existing methods fine-tune pre-trained LLMs on time series tasks and achieves promising results [66, 67, 68], where textual data (e.g., domain instructions, metadata, summaries) are often encoded as prefix embeddings to enrich time series representations [69, 70, 71, 72, 73, 74]. These techniques also contribute to the emergence of time series foundation models [68, 75, 76, 77, 78]. An alternative line of research leverages the zero-shot or few-shot capabilities of LLMs by directly prompting pre-trained language models with text-converted time series [79] or context-laden prompts representing domain knowledge [30, 32, 33], often yielding surprisingly strong performance in real-world scenarios. Furthermore, LLMs can act as knowledge inference modules, synthesizing high-level patterns or explanations that augment standard time series pipelines [80, 31, 81, 82]. Building on this trend, recent works have explored time series reasoning with LLMs by framing tasks as natural-language problems, including time series understanding & question answering [83, 84, 85, 78, 86, 87, 88, 89], and language-guided inference [83, 90, 78, 85]. Compared with existing works, our work provides a unique synergy between an explainable model and LLMs, enhancing explainable prediction through iterative, grounded, and reflective interaction.
@@ 3.1 Problem Statement
In this paper, we consider a multi-modal time series prediction problem. Each instance is represented by the multi-modal input $(\boldsymbol{x},\boldsymbol{s})$, where $\boldsymbol{x}=(x_{1},x_{2},\cdots,x_{T})\in\mathbb{R}^{N\times T}$ denotes time series data with $N$ variables and $T$ historical time steps, and $\boldsymbol{s}$ denotes the corresponding text data describing the real-world context. The text data $\boldsymbol{s}$ can be further divided into $L$ meaningful segments. Based on the historical time series and textual context, our objective is to predict the future outcome $\boldsymbol{y}$, either as a discrete value for classification tasks, or as a continuous value for regression tasks. We focus on the classification task to reflect discrete decision-making scenarios common in real-world multi-modal applications, where interpretability is often essential for reliable decision support. We also include an evaluation of the regression setting in the Appendix F. There are four major components in the proposed TimeXL framework, a multi-modal prototype encoder $\mathcal{M}_{\mathrm{enc}}$ that provides an initial prediction and case-based explanations, a prediction LLM $\mathcal{M}_{\mathrm{pred}}$ that provides a prediction based on contextual understanding with explanations, a reflection LLM $\mathcal{M}_{\mathrm{refl}}$ that generates feedback, and a refinement LLM $\mathcal{M}_{\mathrm{refine}}$ that refines the textual context based on the feedback. Below, we introduce each component and how they synergize toward better prediction and explanation.
Figure 2: The training protocol of TimeXL (left), and multi-modal prototype-based encoder (right).
@@ 3.2.1 Multi-modal Sequence Modeling with Prototypes
Sequence Encoder. To capture both temporal and semantic dependencies, we adopt separate encoders for time series ($\mathcal{E}_{\theta}$) and text ($\mathcal{E}_{\phi}$). For $\boldsymbol{x}\in\mathbb{R}^{N\times T}$, the time series encoder $\mathcal{E}_{\theta}$ maps the entire sequence into one or multiple representations, which serve as candidates for prototype learning. Simultaneously, the text input $\boldsymbol{s}$ is first transformed by a frozen pre-trained language model, $\mathrm{PLM}$ (e.g., BERT [91] or Sentence-BERT [92]), to produce embeddings $\boldsymbol{e}_{s}\in\mathbb{R}^{d_{s}\times L}$. These embeddings are then processed by a separate encoder $\mathcal{E}_{\phi}$ to extract meaningful text features. It is worth noting that the choice of $\mathcal{E}_{\theta}$ and $\mathcal{E}_{\phi}$ also affects the granularity of explanations. As we will introduce shortly, the prototypes are learned based on sequence representations and are associated with the counterparts in the input space, where the correspondences are determined by the encoders. In this paper, we choose convolution-based encoders for both modalities to capture the fine-grained sub-sequence (i.e., segment) patterns:
$\boldsymbol{Z}_{\mathrm{time}}=\bigl{(}\boldsymbol{z}_{1},\dots,\boldsymbol{z}_{T-w+1}\bigr{)}\;=\;\mathcal{E}_{\theta}(\boldsymbol{x}),\quad\boldsymbol{Z}_{\mathrm{text}}=\bigl{(}\boldsymbol{z}^{\prime}_{1},\dots,\boldsymbol{z}^{\prime}_{L-w^{\prime}+1}\bigr{)}\;=\;\mathcal{E}_{\phi}(\boldsymbol{e}_{s})$ (1)
@@
Prototype Allocation. To establish interpretability, we learn a set of time series prototypes and text prototypes for each class $c\in\{1,\dots,C\}$. Specifically, we introduce: $\boldsymbol{P}^{(c)}_{\mathrm{time}}\in\mathbb{R}^{k\times h},\boldsymbol{P}^{(c)}_{\mathrm{text}}\in\mathbb{R}^{k^{\prime}\times h^{\prime}}$ so that each prototype $\boldsymbol{p}_{i}^{(c)}\in\mathbb{R}^{h}$ (time series) or $\boldsymbol{p}_{i}^{\prime(c)}\in\mathbb{R}^{h^{\prime}}$ (text) resides in the same feature space as the relevant encoder outputs. For an input sequence, we measure the similarity between each prototype and the most relevant segment in the corresponding modality:
$\mathrm{Sim}_{i}^{(c)}=\operatorname{max}(\mathrm{Sim}_{i,1}^{(c)},\cdots,\mathrm{Sim}_{i,T-w+1}^{(c)}),\text{where}\;\;\mathrm{Sim}_{i,j}^{(c)}=\exp\left(-\left\|\boldsymbol{p}^{(c)}_{i}-\boldsymbol{z}_{j}\right\|_{2}^{2}\right)\in[0,1]$ (2)
@@ 3.2.2 Learning Prototypes toward Better Explanation
$\mathcal{L}_{c}=\sum_{\boldsymbol{z}_{j}\in\boldsymbol{Z}_{(\cdot)}}\min_{\boldsymbol{p}_{i}\in\boldsymbol{P}_{(\cdot)}}\left\|\boldsymbol{z}_{j}-\boldsymbol{p}_{i}\right\|_{2}^{2},\quad\mathcal{L}_{e}=\sum_{\boldsymbol{p}_{i}\in\boldsymbol{P}_{(\cdot)}}\min_{\boldsymbol{z}_{j}\in\boldsymbol{Z}_{(\cdot)}}\left\|\boldsymbol{p}_{i}-\boldsymbol{z}_{j}\right\|_{2}^{2}$ (4)
@@ 3.3.1 Model Synergy for Augmented Prediction
Prediction with Enriched Contexts. The prediction LLM agent $\mathcal{M}_{\mathrm{pred}}$ generates predictions based on the input text $\boldsymbol{s}$. To improve prediction accuracy, the encoder $\mathcal{M}_{\mathrm{enc}}$ supplements $\boldsymbol{s}$ with case-based explanations. Specifically, $\mathcal{M}_{\mathrm{enc}}$ selects the $\omega$ prototypes that exhibit the highest relevance to any of the textual segments within $\boldsymbol{s}$. Relevance is determined by the similarity scores used in Equation 2. These selected prototypes are then added to the input prompt of $\mathcal{M}_{\mathrm{pred}}$ as explanations, providing richer real-world context and leading to more accurate predictions. The $\omega$ prototype-segment pairs, which construct the explanation $\mathbf{expl}_{s}$ of the input text $\boldsymbol{s}$, are retrieved as follows:
$\mathbf{expl}_{s}=\left\{\left(\boldsymbol{p}_{i}^{(c)},\boldsymbol{s}_{j}\right):(i,j,c)\in\operatorname{Top-\omega}(\mathrm{Sim}_{\mathrm{text}})\right\},\operatorname{Top-\omega}(\mathrm{Sim}_{\mathrm{text}})=\operatorname{argTop-\omega}_{(i,j,c)}\left(\mathrm{Sim^{\prime}}_{i,j}^{(c)}\right)$
Note that $i,j,c$ denotes the prototype index, segment index, and class index, respectively. As $\mathbf{expl}_{\boldsymbol{s}}$ can contain relevant contextual guidance across multiple classes, it augments the input space and removes semantic ambiguity for prediction agent $\mathcal{M}_{\mathrm{pred}}$. Therefore, the prediction is drawn as $\boldsymbol{\hat{y}}_{\mathrm{LLM}}=\mathcal{M}_{\mathrm{pred}}(\boldsymbol{s},\mathbf{expl}_{\boldsymbol{s}})$. The prompt for querying $\mathcal{M}_{\mathrm{pred}}$ is provided in Appendix D, Figure 14.
Fused Predictions. We compile the final prediction based on a fusion of both the multi-modal encoder $\mathcal{M}_{\mathrm{enc}}$ and prediction LLM $\mathcal{M}_{\mathrm{pred}}$. Specifically, we linearly combine the continuous prediction probabilities $\boldsymbol{\hat{y}}_{\mathrm{enc}}$ and discrete prediction $\boldsymbol{\hat{y}}_{\mathrm{LLM}}$: $\boldsymbol{\hat{y}}=\alpha\boldsymbol{\hat{y}}_{\mathrm{enc}}+(1-\alpha)\boldsymbol{\hat{y}}_{\mathrm{LLM}}$, where $\alpha\in[0,1]$ is the hyperparameter selected from validation data. The encoder $\mathcal{M}_{\mathrm{enc}}$ and prediction agent $\mathcal{M}_{\mathrm{pred}}$ enhance each other based on their unique strengths. The $\mathcal{M}_{\mathrm{enc}}$ is fine-tuned based on explicit supervised signals, ensuring accuracy in capturing temporal and contextual dependencies of multi-modal time series. On the other hand, $\mathcal{M}_{\mathrm{pred}}$ contributes deep semantic understanding drawn from extensive text corpora. By fusing predictions from two distinct perspectives, we achieve a synergistic augmentation toward more accurate predictions for complex multi-modal time series.